# 3D and 4D World Modeling: A Survey

### v2

@@ Abstract
World modeling has become a cornerstone in AI research, enabling agents to understand, represent, and predict the dynamic environments they inhabit. While prior work largely emphasizes generative methods for 2D image and video data, they overlook the rapidly growing body of work that leverages native 3D and 4D representations such as RGB-D imagery, occupancy grids, and LiDAR point clouds for large-scale scene modeling. At the same time, the absence of a standardized definition and taxonomy for “world models” has led to fragmented and sometimes inconsistent claims in the literature. This survey addresses these gaps by presenting the first comprehensive review explicitly dedicated to 3D and 4D world modeling and generation. We establish precise definitions, introduce a structured taxonomy spanning video-based (VideoGen), occupancy-based (OccGen), and LiDAR-based (LiDARGen) approaches, and systematically summarize datasets and evaluation metrics tailored to 3D/4D settings. We further discuss practical applications, identify open challenges, and highlight promising research directions, aiming to provide a coherent and foundational reference for advancing the field. A systematic summary of existing literature is available at https://github.com/worldbench/survey.
Index Terms:
World Models; Video Generation; 3D Generation; 4D Generation; 3D Scene Understanding; Spatial Intelligence
Figure 1: Outline of the survey. This work focuses on native 3D and 4D representations: video streams, occupancy grids, and LiDAR point clouds, guided by geometric ($\mathcal{C}_{\mathrm{geo}}$), action-based ($\mathcal{C}_{\mathrm{act}}$), and semantic ($\mathcal{C}_{\mathrm{sem}}$) conditions (Sec. 2). Methods are framed under two paradigms, generative (synthesis from observations and conditions) and predictive (forecasting from history and actions), and grouped into four functional types (Sec. 3). We cover three modality tracks and standardize evaluations (Sec. 4), applications (Sec. 5), and future trends (Sec. 6) across diverse generation, forecasting, and downstream tasks.
@@ 1 Introduction
Beyond these native formats, world modeling has also been explored in adjacent domains [27, 28, 29]. Some works address video, panoramic, or mesh-based data, with systems of this kind providing large-scale, general-purpose video-mesh generation capabilities [30, 31]. In parallel, another line of research focuses on 3D object generation for asset creation, which specializes in controllable and high-fidelity object synthesis [32, 33, 34]. Meanwhile, industrial projects from leading companies have launched ambitious world modeling initiatives that target practical applications ranging from interactive robotics and immersive simulation to large-scale digital twins [35, 36, 37, 38, 39, 40], underscoring the growing importance of this field in both academia and industry.
@@
  • Why native 3D and 4D matters? Unlike 2D projections, native 3D/4D signals encode metric geometry, visibility, and motion in the coordinates where physics acts [49, 18]. This makes them first-class carriers of constraints needed for actionable modeling: multi-view and egocentric consistency, rigid-body and non-rigid kinematics, scene-scale occlusion reasoning, and map/topology adherence. In safety-critical settings, agents must not only produce photorealistic frames but also obey geometry, causality, and controllability; RGB-D, occupancy, and LiDAR provide the inductive bias to satisfy these requirements. Sec. 2 will formalize these representations and the conditioning signals ($\mathcal{C}_{\mathrm{geo}},\mathcal{C}_{\mathrm{act}},\mathcal{C}_{\mathrm{sem}}$) we use throughout the survey.
  • Position in the broader landscape. The adjacent lines – video/panorama/mesh world models [30, 31] and object-centric 3D asset generators [32, 33, 34] – are complementary: they supply appearance, topology, and assets, while native 3D/4D world models supply geometry-grounded dynamics and interaction [19, 50]. Practical systems increasingly compose these capabilities: mesh/panorama worlds initialized from assets, then driven by occupancy- or LiDAR-based dynamics, or video models constrained by 3D priors for view and motion correctness. Our scope centers on the latter – native 3D/4D – while acknowledging and cross-referencing where cross-pollination occurs.
  • From conditions to functions. A common pain point in the field is conflating “what the model consumes” (conditions) with “what the model does” (function). We therefore separate the roles of geometry/action/semantics conditions (Table I) from functional types. Sec. 3 organizes methods by representation modality –VideoGen, OccGen, LiDARGen – and then by four functional roles: 1Data Engines (diverse scene synthesis under $\mathcal{C}_{\mathrm{geo}},\mathcal{C}_{\mathrm{sem}},\mathcal{C}_{\mathrm{act}}$), 2Action Interpreters (forecasting under $\mathcal{C}_{\mathrm{act}}$ with history), 3Neural Simulators (closed-loop rollouts with policy-in-the-loop), and 4Scene Reconstructors (completion/retargeting from partial 3D/4D observations). This decoupling lets us compare heterogeneous methods on common axes of fidelity, consistency, controllability, and scalability.
@@
Figure 2: Summary of representative video-based generation (VideoGen), occupancy-based generation (OccGen), and LiDAR-based generation (LiDARGen) models from existing literature. For the complete list of related methods and discussions on their specifications, configurations, and technical details, kindly refer to Sec. 3.1, Sec. 3.2, and Sec. 3.3, respectively.
Figure 3: Summary of existing datasets and benchmarks used for training and evaluating VideoGen, OccGen, and LiDARGen models. For detailed dataset configurations and statistics, kindly refer to Table V. Images adopted from the original papers.
@@ 2.2 Definition of World Modeling in 3D and 4D
The above scene representations form the structural backbone of 3D/4D world models. In practice, generating or forecasting them requires additional conditions – auxiliary signals that constrain spatial structure, describe agent behavior, or define high-level semantics. As summarized in Table I, these conditions are typically grouped into:
@@
2.2.1 Model Definitions
Predictive World Models instead aim to forecast the future evolution of the scene based on historical observations, often under action conditions that describe planned or executed agent behavior. This process can be formulated as:
TABLE I: Summary of the rich collection of conditions used by existing VideoGen, OccGen, and LiDARGen models. The conditions are categorized into three main groups: geometric conditions, action-based conditions, and semantic conditions. The tasks are [Uncaptioned image] video generation (Sec. 3.1), [Uncaptioned image] occupancy generation (Sec. 3.2), and [Uncaptioned image] LiDAR generation (Sec. 3.3).
@@ 3.1 World Modeling from Video Generation
Video-based generation has emerged as a new paradigm, offering visual cues and temporal dynamics to model complex real-world scenarios. By generating multi-view or egocentric video sequences, these models can synthesize training data, predict future outcomes, and create interactive simulation environments. Based on their primary function, existing methods can be grouped into three categories: [Uncaptioned image] Data Engines, [Uncaptioned image] Action Interpreters, and [Uncaptioned image] Neural Simulators. Table II summarizes existing models under these domains.
@@ 3.1.1 Data Engines
Perception Data Augmentation. Generative scene synthesis alleviates real-world data scarcity and addresses long-tail perception challenges. Early work focused on BEV-guided realistic street scenes. BEVGen [72] uses an autoregressive transformer and cross-view transformation to produce spatially consistent surrounding images aligned with a given BEV layout. BEVControl [73] centers on diffusion models to boost the quality of synthetic data, particularly for augmenting challenging long-tail scenarios. Subsequently, MagicDrive [20] made significant progress in driving scene generation and data augmentation, combining 3D geometry and semantic descriptions, and camera poses to generate high-fidelity images. Later work introduced finer conditioning. For instance, SyntheOcc [75] uses 3D semantic multi-plane images for comprehensive, spatially aligned conditioning, and PerLDiff [79] proposes perspective-layout diffusion models that fully leverage perspective 3D geometry to enhance realism and consistency. On the other hand, approaches such as Panacea [80], DrivingDiffusion [81], and SubjectDrive [82] introduce 4D attention, keyframes, and subject control to improve the temporal consistency and data diversity of 3D controllable multi-view videos. NoiseController [78] proposes multi-level noise decomposition and multi-frame collaborative denoising to enhance spatiotemporal coherence. For long-horizon video generation, DiVE [76], MagicDrive-V2 [74], and Cosmos-Drive [35] leverage the flexibility and scalability of DiT to produce longer videos. Glad [83] uses latent-variable propagation, and STAGE [84] uses hierarchical temporal feature transfer to generate long videos in a streaming fashion. Others like UniScene [77] and BEVWorld [85] explore multi-modal data synthesis to broaden applications, supporting downstream perception tasks that leverage information from multiple modalities. These advances enable robust, scalable autonomous driving perception systems by delivering diverse, controllable, and long-horizon training data that capture real-world variability.
@@
TABLE II: Summary of video-based generation (VideoGen) models.
@@ B   C   T
S 12Hz SD VG, 3SR $Uncaptioned image] \faExternalLink$
@@
S 2Hz MLLM VG, E2E $Uncaptioned image] \faExternalLink$
@@ H   T   C   T
S - SD VG, 3SR $Uncaptioned image] \faExternalLink$
@@ H   C   T
L 12Hz SD, DiT VG $Uncaptioned image] \faExternalLink$
@@ B   T   C
[…]
@@ H   C   T   P
[…]
@@
S 2Hz DiT VG $Uncaptioned image] \faExternalLink$
@@ B   C   T
L 12Hz DiT VG, 3SR $Uncaptioned image] \faExternalLink$
@@
[…]
3.1.2 Action Interpreters
Action-Guided Video Generation. Action-conditioned generation models empower agents to predict future outcomes based on intended maneuvers, effectively bridging low-level control inputs with high-fidelity video rollouts of plausible futures. GAIA-1 [60] pioneered a generative model that fuses video, text, and action inputs to synthesize realistic driving scenarios with detailed control over ego-vehicle behavior and scene attributes. GAIA-2 [126] expanded this framework to include agent configurations, environmental factors, and road semantics. GenAD [96] further enhanced generalization by releasing the OpenDV dataset alongside a predictive model that supports zero-shot, language- and action-conditioned predictions. Vista [103] applies robust action conditioning across diverse scenarios, while GEM [116] delivers multimodal outputs with precise ego-motion control, and MaskGWM [117] boosts fidelity and long-horizon predictions using mask-based diffusion. To address error accumulation in long video synthesis, InfinityDrive [107] and Epona [59] proposed memory injection and a chain-of-forward training strategy, respectively. In addition, DrivingWorld [112] generates scenarios from predefined trajectories, functioning as a neural driving simulator. Other approaches, such as DriVerse [128], MiLA [125], PosePilot [129], and LongDWM [131], focus on trajectory alignment, temporal stability, pose controllability, and depth-free guidance. Collectively, these advances drive action-conditioned generation toward better precision, temporal coherence, and robustness.
Forecasting-Driven Action Planning. Another line of work forecasts future states from current observations and ego actions, letting planners evaluate outcomes before committing [132, 133, 134]. Different from purely reactive schemes, these approaches emphasize anticipatory decision-making, allowing the agent to virtually “test” multiple futures and avoid unsafe trial-and-error in the real world. Drive-WM [102] generates video rollouts of candidate maneuvers, scoring them with image-based rewards for trajectory selection. DriveDreamer [58] proposed the ActionFormer to predict future states and ego-environment interactions. ADriver-I [101] combines multimodal LLMs with autoregressive control signals and world evolution prediction. Vista [103] incorporates uncertainty-aware reward modules for robust action evaluation. GPT-style designs such as DrivingGPT [111] and DrivingWorld [112] model visual and action tokens jointly for planning via next-token prediction. Integrated frameworks like Doe-1 [109] unify perception, prediction, and planning for closed-loop autonomous driving, while VaVAM [122] bridges video diffusion and an action expert for decision-making. ProphetDWM [130] further couples latent action learning with state forecasting for long-term planning. Overall, by simulating diverse futures and leveraging feedback, forecast-driven models enhance generalization and safety in end-to-end autonomous driving.
@@ 3.1.3 Neural Simulators
TABLE III: Summary of occupancy-based generation (OccGen) models.
@@
$\bullet$ Categories: [Uncaptioned image] Scene Representor (Sec. 3.2.1), [Uncaptioned image] Occ Forecaster (Sec. 3.2.2), and [Uncaptioned image] Autoregressive Simulator (Sec. 3.2.3).
@@ 11 Emergent-Occ [167] ECCV’22
[…]
@@ 22 FF4D [63] CVPR’23
[…]
@@ 33 SSD [168] arXiv’23
[…]
@@ 44 UniWorld [169] arXiv’23
[…]
@@ 55 UniScene [170] RA-L’24
[…]

### v3

# ] [Uncaptioned image] WorldBench Team

@@ Abstract
World modeling has become a cornerstone in AI research, enabling agents to understand, represent, and predict the dynamic environments they inhabit. While prior work largely emphasizes generative methods for 2D image and video data, they overlook the rapidly growing body of work that leverages native 3D and 4D representations such as RGB-D imagery, occupancy grids, and LiDAR point clouds for large-scale scene modeling. At the same time, the absence of a standardized definition and taxonomy for “world models” has led to fragmented and sometimes inconsistent claims in the literature. This survey addresses these gaps by presenting the first comprehensive review explicitly dedicated to 3D and 4D world modeling and generation. We establish precise definitions, introduce a structured taxonomy spanning video-based (VideoGen), occupancy-based (OccGen), and LiDAR-based (LiDARGen) approaches, and systematically summarize datasets and evaluation metrics tailored to 3D/4D settings. We further discuss practical applications, identify open challenges, and highlight promising research directions, aiming to provide a coherent and foundational reference for advancing the field. A systematic summary of existing literature is available on our project page.
Figure 1: Outline of the survey. This work focuses on native 3D and 4D representations: video streams, occupancy grids, and LiDAR point clouds, guided by geometric ($\mathcal{C}_{\mathrm{geo}}$), action-based ($\mathcal{C}_{\mathrm{act}}$), and semantic ($\mathcal{C}_{\mathrm{sem}}$) conditions (Sec. 2). Methods are framed under two paradigms, generative (synthesis from observations and conditions) and predictive (forecasting from history and actions), and grouped into four functional types (Sec. 3). We cover three modality tracks and standardize evaluations (Sec. 4), practical applications (Sec. 5), and future trends (Sec. 6) across diverse generation, forecasting, and downstream task perspectives.
@@ 1 Introduction
Beyond these native formats, world modeling has also been explored in adjacent domains [59, 60, 289]. Some works address video, panoramic, or mesh-based data, with systems of this kind providing large-scale, general-purpose video-mesh generation capabilities. In parallel, another line of research focuses on 3D object generation for asset creation, which specializes in controllable and high-fidelity object synthesis. Meanwhile, industrial projects from leading companies have launched ambitious world modeling initiatives that target practical applications ranging from interactive robotics and immersive simulation to large-scale digital twins [194, 8, 1, 4, 10, 5], underscoring the growing importance of this field in both academia and industry.
@@
Figure 2: Summary of representative video-based generation (VideoGen), occupancy-based generation (OccGen), and LiDAR-based generation (LiDARGen) models from existing literature. For the complete list of related methods and discussions on their specifications, configurations, and technical details, kindly refer to Sec. 3.1, Sec. 3.2, and Sec. $3.$
Why native 3D and 4D matters?
Unlike 2D projections, native 3D/4D signals encode metric geometry, visibility, and motion in coordinates where physics acts [139, 17]. This makes them first-class carriers of constraints needed for actionable modeling: multi-view and egocentric consistency, rigid-body and non-rigid kinematics, scene-scale occlusion reasoning, and map/topology adherence. In safety-critical settings, agents must not only produce photorealistic frames but also obey geometry, causality, and controllability; RGB-D, occupancy, and LiDAR provide the inductive bias to satisfy these requirements. Sec. 2 will formalize these representations and the conditioning signals ($\mathcal{C}_{\mathrm{geo}},\mathcal{C}_{\mathrm{act}},\mathcal{C}_{\mathrm{sem}}$) we use throughout the survey.
Position in the broader landscape.
The adjacent lines – video/panorama/mesh world models [97] and object-centric 3D asset generators [277] – are complementary: they supply appearance, topology, and assets, while native 3D/4D world models supply geometry-grounded dynamics and interaction [296, 304]. Practical systems increasingly compose these capabilities: mesh/panorama worlds initialized from assets, then driven by occupancy- or LiDAR-based dynamics, or video models constrained by 3D priors for view and motion correctness. Our scope centers on the latter – native 3D/4D – while acknowledging and cross-referencing where cross-pollination occurs.
From conditions to functions.
A common pain point in the field is conflating “what the model consumes” (conditions) with “what the model does” (function). We therefore separate the roles of geometry/action/semantics conditions (Table 1) from functional types. Sec. 3 organizes methods by representation modality –VideoGen, OccGen, LiDARGen – and then by four functional roles: 1Data Engines (diverse scene synthesis under $\mathcal{C}_{\mathrm{geo}},\mathcal{C}_{\mathrm{sem}},\mathcal{C}_{\mathrm{act}}$), 2Action Interpreters (forecasting under $\mathcal{C}_{\mathrm{act}}$ with history), 3Neural Simulators (closed-loop rollouts with policy-in-the-loop), and 4Scene Reconstructors (completion/retargeting from partial 3D/4D observations). This decoupling lets us compare heterogeneous methods on common axes of fidelity, consistency, controllability, and scalability.
@@ 2 Preliminaries
Figure 3: Summary of existing datasets & benchmarks used for training and evaluating VideoGen, OccGen, and LiDARGen models. For detailed configurations and statistics, kindly refer to Table 5. Images adopted from the original papers.
@@ 2.2 Definition of World Modeling in 3D and 4D
The above scene representations form the structural backbone of 3D/4D world models. In practice, generating or forecasting them requires additional conditions – auxiliary signals that constrain spatial structure, describe agent behavior, or define high-level semantics. As summarized in Table 1, these conditions are typically grouped into:
@@
Table 1: Summary of the rich collection of conditions used by VideoGen, OccGen, and LiDARGen models. The conditions are categorized into three main groups: geometric conditions, action-based conditions, and semantic conditions. The tasks are [Uncaptioned image] video generation (Sec. 3.1), [Uncaptioned image] occupancy generation (Sec. 3.2), and [Uncaptioned image] LiDAR generation (Sec. 3.3).
@@
2.2.1 Model Definitions
Predictive World Models instead aim to forecast the future evolution of the scene based on historical observations, often under action conditions that describe planned or executed agent behavior. This process is formulated as:
@@ 3.1 World Modeling from Video Generation
Video-based generation has emerged as a new paradigm, offering visual cues and temporal dynamics to model complex real-world scenarios. By generating multi-view or egocentric video sequences, these models can synthesize training data, predict future outcomes, and create interactive simulation environments. Based on their primary function, existing methods can be grouped into three categories: [Uncaptioned image] Data Engines, [Uncaptioned image] Action Interpreters, and [Uncaptioned image] Neural Simulators. Table 2 summarizes existing models under these domains.
@@ 3.1.1 Data Engines
Perception Data Augmentation. Generative scene synthesis alleviates real-world data scarcity and addresses long-tail perception challenges. Early work focused on BEV-guided realistic street scenes. BEVGen [211] uses an autoregressive transformer and cross-view transformation to produce spatially consistent surrounding images aligned with a given BEV layout. BEVControl [275] centers on diffusion models to boost the quality of synthetic data, particularly for augmenting challenging long-tail scenarios. Subsequently, MagicDrive [64] made significant progress in driving scene generation and data augmentation, combining 3D geometry and semantic descriptions, and camera poses to generate high-fidelity images. Later work introduced finer conditioning. For instance, SyntheOcc [131] uses 3D semantic multi-plane images for comprehensive, spatially aligned conditioning, and PerLDiff [293] proposes perspective-layout diffusion models that fully leverage perspective 3D geometry to enhance realism and consistency. On the other hand, approaches such as Panacea [244], DrivingDiffusion [135], and SubjectDrive [93] introduce 4D attention, keyframes, and subject control to improve the temporal consistency and data diversity of 3D controllable multi-view videos. NoiseController [47] proposes multi-level noise decomposition and multi-frame collaborative denoising to enhance spatiotemporal coherence.
For long-horizon video generation, DiVE [104], MagicDrive-V2 [63], and Cosmos-Drive [194] leverage the flexibility and scalability of DiT to produce longer videos. Glad [256] uses latent-variable propagation, and STAGE [228] uses hierarchical temporal feature transfer to generate long videos in a streaming fashion. Others like UniScene [125] and BEVWorld [299] explore multi-modal data synthesis to broaden applications, supporting downstream perception tasks that leverage information from multiple modalities. These advances enable robust, scalable autonomous driving perception systems by delivering diverse, controllable, and long-horizon training data that capture real-world variability.
@@
3.1.2 Action Interpreters
Action-Guided Video Generation. Action-conditioned generation models empower agents to predict future outcomes based on intended maneuvers, effectively bridging low-level control inputs with high-fidelity video rollouts of plausible futures. GAIA-1 [86] pioneered a generative model that fuses video, text, and action inputs to synthesize realistic driving scenarios with detailed control over ego-vehicle behavior and scene attributes. GAIA-2 [198] expanded this framework to include agent configurations, environmental factors, and road semantics. GenAD [274] further enhanced generalization by releasing the OpenDV dataset alongside a predictive model that supports zero-shot, language- and action-conditioned predictions. Vista [65] applies robust action conditioning across diverse scenarios, while GEM [80] delivers multimodal outputs with precise ego-motion control, and MaskGWM [178] boosts fidelity and long-horizon predictions using mask-based diffusion.
To address error accumulation in long video synthesis, InfinityDrive [77] and Epona [295] proposed memory injection and a chain-of-forward training strategy, respectively. In addition, DrivingWorld [91] generates scenarios from predefined trajectories, functioning as a neural driving simulator. Other approaches, such as DriVerse [134], MiLA [226], PosePilot [107], and LongDWM [234], focus on trajectory alignment, temporal stability, pose controllability, and depth-free guidance. Collectively, these advances drive action-conditioned generation toward better precision, temporal coherence, and robustness.
Forecasting-Driven Action Planning. Another line of work forecasts future states from current observations and ego actions, letting planners evaluate outcomes before committing [227, 132, 311]. Different from purely reactive schemes, these approaches emphasize anticipatory decision-making, allowing the agent to virtually “test” multiple futures and avoid unsafe trial-and-error in the real world. Drive-WM [239] generates video rollouts of candidate maneuvers, scoring them with image-based rewards for trajectory selection. DriveDreamer [236] proposed the ActionFormer to predict future states and ego-environment interactions. ADriver-I [102] combines multimodal LLMs with autoregressive control signals and world evolution prediction. Vista [65] incorporates uncertainty-aware reward modules for robust action evaluation.
GPT-style designs such as DrivingGPT [35] and DrivingWorld [91] model visual and action tokens jointly for planning via next-token prediction. Integrated frameworks like Doe-1 [309] unify perception, prediction, and planning for closed-loop autonomous driving, while VaVAM [12] bridges video diffusion and an action expert for decision-making. ProphetDWM [233] further couples latent action learning with state forecasting for long-term planning. Overall, by simulating diverse futures and leveraging feedback, forecast-driven models enhance generalization and safety in end-to-end autonomous driving.
Table 2: Summary of video-based generation (VideoGen) models.
@@ B C T
S 12Hz SD VG, 3SR [Uncaptioned image] [Uncaptioned image] W
@@
S 2Hz MLLM VG, E2E [Uncaptioned image] [Uncaptioned image] W
@@ H T C T
S - SD VG, 3SR [Uncaptioned image] [Uncaptioned image] W
@@ H C T
L 12Hz SD, DiT VG [Uncaptioned image] [Uncaptioned image] W
@@ B T C
[Uncaptioned image] [Uncaptioned image] N/A L 2Hz MLLM VG [Uncaptioned image] W
@@ H C T P
[Uncaptioned image] [Uncaptioned image] N/A S 12Hz SD VG, 3SR [Uncaptioned image] W
@@
S 2Hz DiT VG [Uncaptioned image] [Uncaptioned image] W
@@ B C T
L 12Hz DiT VG, 3SR [Uncaptioned image] [Uncaptioned image] W
@@
L 12Hz SD VG [Uncaptioned image] W
@@ 3.1.3 Neural Simulators
Table 3: Summary of occupancy-based generation (OccGen) models.
@@
$\bullet$ Categories: [Uncaptioned image] Scene Representor (Sec. 3.2.1), [Uncaptioned image] Occ Forecaster (Sec. 3.2.2), and [Uncaptioned image] AR Simulator (Sec. 3.2.3).
@@ 11 Emergent-Occ [112] ECCV’22
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] N/A 7 Enc-Dec OF, E2E [Uncaptioned image] W
@@ 22 FF4D [113] CVPR’23
[Uncaptioned image] [Uncaptioned image] N/A 5 Enc-Dec OF [Uncaptioned image] W
@@ 33 SSD [123] arXiv’23
[Uncaptioned image] [Uncaptioned image] N/A 1 LDM O3G [Uncaptioned image] W
@@ 44 UniWorld [170] arXiv’23
[Uncaptioned image] [Uncaptioned image] N/A - Enc-Dec PT [Uncaptioned image] W
@@ 55 UniScene [168] RA-L’24
[Uncaptioned image] [Uncaptioned image] N/A - Enc-Dec PT [Uncaptioned image] W